OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

apply the same steps I showed you earlier. it follows a similar pattern. can you identify the first step?

OpenStudy (id21):

-2x-6=-2x-2-4 -6=-6 x can be any number

OpenStudy (anonymous):

-2x-6 = -2x-2-4 -2x-2x=-6+6 0=0

OpenStudy (anonymous):

so its no soulution? or all real numbers?

OpenStudy (id21):

All reals

OpenStudy (anonymous):

-2x-6=-2x-2-4 -6=-6 0=0 it means all real numbers

OpenStudy (anonymous):

okay thank you